# Generalized Landau Yang Theorem

### v1

@@ 5.1 Stueckelberg extension of SUSY
Lastly, we can extend the above ideas and consider the Stueckelberg extension of supersymmetric QED interacting with a massive vector supermultiplet by providing mass to the photon and its superpartner the photino. For this we need a chiral multiplet in addition to the vector multiplet (for the photon) to cancel the supersymmetric gauge transformations. The chiral multiplet is the supersymmetric analogue of the Stueckelberg field. We refer to the papers for such an extension[9, 18, 19].
@@ 6 Conclusions
Landau Yang theorem is an important selection rule which has helped in understanding the role of symmetries in Particle Physics. Since it is related to the massless particles it has acquired importance coming from asymptotic properties of fields. It has also made the understanding of the Hilbert space of states with appropriate properties the central focus. In this connection the mass of the particle as a Lorentz covariant regulator for infrared questions is crucial. Similarly semi-inclusive cross sections for S-matrix in scalar QED, for instance, seem to require mass for the photon as a regulator to remove ambiguities in handling infrared divergences [20]. Furthermore the Landau Yang theorem can also be extended to identical massless fermions. In addition we have also provided extension of the theorem to supersymmetry in this paper.
As Schrodinger[13] pointed out it is important to understand the approach to masslessness of particles even though nature may require zero mass.
@@
A small mass may provide some novel ways of understanding some cosmological questions on the nature of dark matter/energy[21]. Higgs mechanism through which photon can acquire mass may also provide a link between primordial blackholes and dark matter in early galaxies [22, 23]. It is expected that at length scales $L~{}\leq~{}\frac{h}{m_{\gamma}}$ the longitudinal photon behaves like a Stueckelberg scalar. But experimental observation of emission of such a scalar is very difficult if the mass is small as explained by Schrodinger[13]. The only observational signature would be missing energy and momentum in the process $Z\longrightarrow\gamma_{T}~{}+~{}\gamma_{L}$ (which could be a fraction of the energy-momentum carried by the transverse photon, although the rate of such processes will be limited by $M_{Z}(m_{\gamma}/m_{e})^{2}$, as seen earlier).
Acknowledgements: TRG would like to thank Alok Laddha and K. S. Narayan for discussions. Feynman diagrams were made using Feynman Diagram Maker (https://www.aidansean.com/feynman/).

### v2

@@ 5.1 Stueckelberg extension of SUSY
Lastly, we can extend the above ideas and consider the Stueckelberg extension of supersymmetric QED interacting with a massive vector supermultiplet by providing mass to the photon and its superpartner the photino. For this we need a chiral multiplet in addition to the vector multiplet (for the photon) to cancel the supersymmetric gauge transformations. The chiral multiplet is the supersymmetric analogue of the Stueckelberg field. We refer to the literature for such an extension buchbinder ; nath ; trg .
@@ 6 Conclusions
Landau Yang theorem is an important selection rule which has helped in understanding the role of symmetries in Particle Physics. While the original work landau ; yang showed the impossibility of the decay of a massive spin one particle to two identical massless bose particles, the conclusion is true even for identical massless fermions of the same helicity bal ; Bal-Book . In the present work we have also provided extension of the theorem to supersymmetry which, very interestingly, prohibits the decay of a massive spin one-half particle to even two non-identical massless particles which, nevertheless, are part of the same supersymmetric multiplet. In particular, we showed that in the supersymmetric extension of the SM, the zino cannot decay to a photon and a photino.
As the Landau Yang theorem is related to massless particles it has acquired importance coming from asymptotic properties of fields. It has also made the understanding of the Hilbert space of states with appropriate properties the central focus. In this connection the mass of the particle as a Lorentz covariant regulator for infrared questions is crucial. Similarly semi-inclusive cross sections for S-matrix in scalar QED, for instance, seem to require mass for the photon as a regulator to remove ambiguities in handling infrared divergences laddha . As seen, the theorem can be evaded when the photon has a nonzero mass. This is achieved without violating the Lorentz invariance by using the Stueckelberg mechanism as opposed to some of the previous studies where the theorem could be evaded only at the cost of Lorentz invariance rindani ; qcd ; balnew .
It is easy to understand the role of mass as a regulator for this theorem. The process $Z\rightarrow 2\gamma$ can come from a gauge invariant term of the form $S_{int}~{}=\int d^{4}x~{}g(F_{\mu\rho}F_{\nu\sigma}g^{\rho\sigma})(\partial^{\mu}~{}V^{\nu})$ where $V^{\mu}$ is the massive spin one vector field and $F_{\mu\nu}$ is the electromagnetic field strength. Integrating by parts leads to $\partial_{\mu}~{}F^{\mu\nu}$ which is zero onshell. If the photon has a small mass, it will provide a nonzero result. The same is true even for supersymmetric theories as well where there will be an additional term in $S_{int}$ given by $\psi^{T}[\gamma_{\mu},\gamma_{\nu}]\psi\partial^{\mu}V^{\nu}$ corresponding to the photino. Once again, in the massless theory this contribution will be zero on-shell as can be confirmed by integrating by parts implying that $Z\rightarrow\tilde{\gamma}\tilde{\gamma}$ is also not allowed. This is consistent with what has been said earlier regarding the decay of a massive spin one particle to two Majorana fermions. But with broken supersymmetry, the photino will become massive and the decay of $Z$ to two photinos becomes allowed. Phenomenologically this is interesting since such decays, if observed, can provide important information about the scale of SUSY.
As Schrodinger schrodinger pointed out it is important to understand the approach to masslessness of particles even though nature may require zero mass.
@@
A small photon mass may provide some novel ways of understanding some cosmological questions on the nature of dark matter/energy trgdark as well. Higgs mechanism through which photon can acquire mass may also provide a link between primordial blackholes and dark matter in early galaxies isra ; SMBH . It is expected that at length scales $L~{}\leq~{}\frac{h}{m_{\gamma}}$ the longitudinal photon behaves like a Stueckelberg scalar. But experimental observation of emission of such a scalar is very difficult if the mass is small as explained by Schrodinger schrodinger . The only observational signature would be missing energy and momentum in the process $Z\longrightarrow\gamma_{T}~{}+~{}\gamma_{L}$ (which could be a fraction of the energy-momentum carried by the transverse photon, although the rate of such processes will be limited by $M_{Z}(m_{\gamma}/m_{e})^{2}$, as seen earlier).
Acknowledgements.
TRG would like to thank Alok Laddha and K. S. Narain for discussions. Authors would like to thank Parmeshwaran Nair for his comments on the draft. Feynman diagrams were made using Feynman Diagram Maker (https://www.aidansean.com/feynman/).